solving systems of equations. solve the system: y=5-3x and 3x+7y=15
Answer by checkley77(7072)

y=5-3x
3x+7y=15 substitute (5-3x) for y in this equation & solve for x.
3x+7(5-3x)=15
3x+35-21x=15
-18x=15-35
-18x=-20
x=-20/-18
x=10/9 ans.
y=5-3*10/9
y=5-30/9
y=(5*9-30)/9
y=(45-30)/9
y=15/9 ans.
Proof
3*10/9+7*15/9=15
30/9+105/9=15
135/9=15
15=15